OpenStudy (anonymous):

Which of the following is a counterexample for the given statement. 6^n*6(n)is divisible by 12. a) n=1 b) n=2 c) n=3 d) n=4

OpenStudy (helder_edwin):

it is \[ \large 6^n\cdot6n \] right?

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

just plug in the numbers.

OpenStudy (helder_edwin):

for instance. if n=3 then \[ \large 6^n\cdot6n=6^3\cdot6\cdot3=216\cdot18=3888 \]

OpenStudy (helder_edwin):

is 3888 divisible by 12??

OpenStudy (anonymous):

right i did that but idk how to tell which ones true

OpenStudy (anonymous):

ohhh i see

OpenStudy (helder_edwin):

use a calculator!!

OpenStudy (anonymous):

i have been

OpenStudy (anonymous):

idk what numbers are divisible by 12 tho

OpenStudy (helder_edwin):

3888 divided by 12 gives u 324 exact. no decimals. so for n=3 it is true

OpenStudy (anonymous):

gotcha thanks!

OpenStudy (helder_edwin):

great